# Understanding while Exploring: Semantics-driven Active Mapping

### v1

@@ 1 Introduction
Mobile robots are expected to play a significant role in human-centered environments, such as warehouses, factories, hospitals, and homes, as well as in dangerous settings, such as mines and nuclear facilities. Rich and accurate geometric and semantic representations are prerequisites in these scenarios so that robots can understand, interpret, and interact meaningfully with their surroundings. For instance, in automated warehouses, robots are required to recognize various items and place them in the correct sorting zones accordingly. Scene understanding is enabled by a semantic map that is linked to the geometric map [1], which represents the spatial layout of an environment, and enriches it with high-level information such as object categories, surface labels, and functional affordances [2, 3, 4, 5]. Such maps are critical for a range of tasks including navigation, inspection, object manipulation, human-robot interaction, and long-term autonomy.
@@
Despite substantial advances in semantic mapping, most current approaches are unable to determine the most informative path for the robot to follow. Instead, they passively rely on externally determined trajectories or predefined exploration strategies [6, 7, 8, 9], often leading to incomplete or suboptimal scene understanding. In this paper, we present an approach for active semantic mapping that seeks to close the loop between perception and action. This allows agents to plan their next moves and observations in order to improve the quality, completeness, and efficiency of the semantic map. Our approach, named ActiveSGM (Active Semantic Gaussian Mapping), is the first active semantic mapping system based on radiance fields, enabling rapid exploration, efficient understanding of the environment, and high-fidelity real-time rendering, ultimately leading to more intelligent and efficient robotic behaviors. ActiveSGM aims to infer the semantic labels of all visible surfaces, without favoring any particular label.
@@
To select the most informative views for the robot, we seek to quantify both geometric and semantic uncertainty. At the geometric level, uncertainty is typically measured by the expected error in the estimated 3D coordinates [10, 11, 12]. At the semantic level, uncertainty estimation primarily captures ambiguity among semantic classes. Recent surveys on semantic uncertainty quantification [13, 14, 15] found that it is inherently dependent upon the choice of semantic representation.
@@
Semantic representation plays a critical role in semantic mapping systems, which commonly adopt two primary forms: probability distributions or embeddings. For distribution-based representations, existing methods (e.g., [16, 17, 18]) employ either hard or soft assignment strategies. Hard assignments, such as one-hot encoding, strictly assign a single label to each pixel. In contrast, soft assignments allocate a complete categorical probability distribution to each 3D primitive, naturally capturing uncertainty but also incurring higher memory requirements as the number of categories grows. Alternatively, embedding-based representations can also be viewed as a form of soft assignment. Methods like  [9, 19] utilize features, such as those from DINO [20] or CLIP [21], to encode semantic embeddings. However, these embeddings are high-dimensional, posing challenges for storage and real-time rendering in large scenes. Consequently, some approaches compress the features into lower-dimensional spaces, such as the three-dimensional RGB color space. The dimension of the embedding feature space directly determines the effectiveness of class discrimination. High-dimensional embeddings, like those from DINO or CLIP, provide a sufficiently expressive feature space to effectively distinguish categories. However, as embeddings become compressed, for instance into RGB space, color blending during multi-view reconstruction inevitably occurs, producing blended colors that may correspond to unrelated categories instead of the original ones.
@@
To summarize, we propose the first dense active semantic mapping system built upon a 3D Gaussian Splatting (3DGS) mapping backbone, which integrates semantic-aware mapping and planning for active reconstruction. This enables the robot to construct a more accurate geometric map and a richer semantic map with fewer observations. Our method addresses several key challenges:
@@ 2 Related Work
In this section, we review prior work, starting from dense SLAM, active mapping, semantic mapping, and concluding with active semantic mapping. We focus on methods utilizing either Neural Radiance Fields (NeRF) [22, 23, 24] or Gaussian Splatting (GS) [25, 26, 27] as the representation.
@@
Autonomous robotics relies on foundational capabilities such as localization, mapping, planning, and motion control [28]. The need to realize these capabilities has spurred advancements in various areas, including visual odometry [29, 30], structure-from-motion (SfM) [31], and Simultaneous Localization and Mapping (SLAM) [32, 33, 34, 10, 35]. For surveys of the impact of radiance fields in SLAM and robotics in general, we refer readers to  [36, 37, 38, 27]. Progress in radiance fields has given rise to a multitude of dense SLAM methods, that estimate depth for almost every pixel of the input images, using NeRF (or other implicit representations, such as TSDF) [39, 40, 41, 42, 43, 44, 45, 46, 47, 48] or GS [49, 50, 51, 52, 53, 54] to represent scene geometry and appearance. We use SplaTAM [51] as the SLAM backbone of our algorithm.
@@
The goal of SLAM is to estimate the camera/vehicle trajectory from sensor data. Active mapping, or exploration, is a related problem in the domain of active perception [55, 56], where the goal is guiding the sensor to acquire images beneficial to a downstream task. The most common objectives are to reduce uncertainty, equivalently to increase information gain, [57] or to detect and visit frontiers [58]. Early work demonstrated the effectiveness of active mapping [59, 60, 61, 62, 63], while overviews of the state of the art can be found in [10, 11, 12].
@@
Recently, NeRF-based approaches have been applied to path planning [64] and next-best-view selection [65, 66, 67], though they are often limited by their high computational cost [68]. To overcome these limitations, hybrid models such as ActiveRMAP [69] integrate implicit and explicit representations.
NARUTO [70] introduces an active neural mapping system with 6DoF movement in unrestricted spaces, while Kuang et al. [68] integrate Voronoi planning to scale exploration to larger environments. 3DGS offers a faster alternative, making real-time mapping and exploration more feasible. Recent works like ActiveSplat [71] utilize a hybrid map with topological abstractions for efficient planning, ActiveGS [72] also uses a hybrid map and associates a confidence with each Gaussian to guide exploration, and AG-SLAM [73] incorporates 3DGS with Fisher Information to balance exploration and localization in complex environments. ActiveGAMER [74] introduces a rendering-based information gain criterion that selects the next-best view for enhancing geometric and photometric reconstruction accuracy in complex environments. RT-GuIDE [75] uses a simple uncertainty measure to achieve real-time planning and exploration on a robot. Recently, NextBestPath [76] considers longer horizons than just the single next view. Like all the methods in this paragraph, however, it does not consider semantics.
@@
The goal of semantic mapping is to infer scene descriptions that go beyond geometry [1]. In general, methods in this category endow their 3D representation with semantic labels, which are inferred via semantic segmentation of the input RGB or RGB-D images. Early work includes approaches such SemanticFusion [77], Fusion++ [4], PanopticFusion [78] and Kimera [6], which have adopted different representations exploring tradeoffs between precision and efficiency. Radiance field-based methods are surveyed by Nguyen et al. [79]. Among them, GSNeRF [80] introduces the Semantic Geo-Reasoning and Depth-Guided Visual modules to train a NeRF that encodes semantics along with appearance. Wilson et al. [15] use the variance of the semantic representation at each Gaussian as a proxy for semantic uncertainty. HUGS [81] jointly optimizes geometry, appearance, semantics, and motion using a combination of static and dynamic 3D Gaussians. Logits for all classes are stored with the Gaussians, but the number of classes is small. All of these approaches operate on all frames in batch mode, however.
@@
Gaussian splats are well suited for semantic mapping because they can encode additional attributes and are amenable to continual learning, unlike NeRF [82]. All methods below operate on RGB-D video inputs. We point out the important representation choices made by their authors. SGS-SLAM [7] augments a GS-based SLAM system with additional test-time supervision via 2D semantic maps. The authors argue that any off-the-self semantic segmentation algorithm can be integrated in SGS-SLAM and use ground truth labels to supervise the splats for simplicity. High-dimensional semantic labels are converted to “semantic colors" to save space. NIDS-SLAM [83] uses a 2D transformer [84] to estimate keyframe semantics, also converting the semantic labels into “semantic colors." NEDS-SLAM [8] reduces the memory footprint of the high-dimensional semantic features obtained by DINO [20] to three values per splat via a lightweight encoder. OpenGS-SLAM [9] infers consistent labels via the consensus of 2D foundational models across multiple views. It can handle an open vocabulary, but stores only one label per splat.
To overcome the limited dimensionality of colormaps, researchers have endowed the splats with embeddings of the high-dimensional vectors of logits. SNI-SLAM [85] model the correlations among appearance, geometry and semantic features through a cross-attention mechanism and use feature planes [41] to save memory. DNS-SLAM [86] relies on a multi-resolution hash-based feature grid. Optimization is performed in latent space, while ground truth 2D semantic maps are used as inputs. SemGauss-SLAM [87] augments the splats with a 16-channel semantic embedding and presents semantic-informed bundle adjustment. The paper includes results using ground truth 2D labels for supervision, as well as labels inferred by a classifier operating on DINOv2 [88] features. Hier-SLAM [5] addresses the increased storage requirements via a hierarchical tree representation, generated by a large language model. It can handle over 500 semantic classes, but it is also provided the ground truth semantic maps of the images during optimization.
@@
All the above approaches are “passive" in the sense that the camera is not actively controlled but follows a predetermined trajectory. Prior relevant research addressed the problem of searching for specific objects leveraging semantic contextual priors, i.e. knowing that cups are typically in the kitchen, but without predicting semantic labels for every point of the map [89, 90, 91]. Among these approaches, more relevant to ours is the work of Zhang et al. [92] that relies on semantic mutual information and properties of the SLAM pose graph for metric-semantic active mapping. An octree is used to maintain the map, but the current implementation is limited to 2D motion and 8 classes, while ground truth labels are used as semantic observations.
Marza et al. [93] added a semantic head to Nerfacto [94] and used it for active mapping of appearance, geometry and semantics. They compared using ground truth semantic labels and Mask-R-CNN [95] to detect 15 object categories, and observed large differences in the metrics. Exploration policies are trained using reinforcement learning and consider the 15 object categories. Unlike our approach, the trajectory is restricted to the ground plane. It is not clear how this approach would have to be modified if all semantic classes in the scene would have to be considered.
@@ 3 Method
Figure 1: Overview of the ActiveSGM System. Our framework integrates observation, mapping, and planning into a unified active semantic mapping system. At each time step, posed RGB-D frames along with semantic predictions from OneFormer [17] are stored in a keyframe database. Selected frames are used to update a Semantic Gaussian Map that encodes geometric, photometric, and semantic properties and is optimized through differentiable rendering. An occupancy-based Exploration Map is updated using the current view and used to sample candidate viewpoints in free space. Next-best views are selected by jointly evaluating geometric and semantic exploration criteria (E.C.), and a path planner navigates toward the selected pose. This closed-loop system enables efficient, semantics-aware reconstruction and exploration in complex 3D environments.
@@ 3.1 Semantic Gaussian Mapping
Gaussian Mapping leverages 3DGS to represent scenes as collections of 3D Gaussians, effectively encoding both appearance and geometry for real-time rendering of high-fidelity color and depth images. Building upon the foundational work of Kerbl et al. [25], we adopt the streamlined approach proposed in SplaTAM [51]. This method employs isotropic Gaussians with view-independent color, optimizing parameters such as color (c), center position ($\boldsymbol{\mu}$), radius ($r$), and opacity ($o$). A notable advantage of 3DGS is its capability for real-time rendering, enabling the synthesis of high-fidelity color and depth images from arbitrary camera poses. This is achieved by transforming 3D Gaussians into camera space, sorting them front-to-back, projecting them onto the 2D image plane, and employing alpha-blending for compositing. The color, depth, and silhouette at pixel $\mathbf{p}$ are rendered from the Gaussian map, where the silhouette indicates whether $\mathbf{p}$ receives a significant projection from any Gaussian. The general rendering process is formulated as
@@
We incorporate semantics into the 3DGS map by using OneFormer [17], a state-of-the-art model for unified segmentation, to perform semantic segmentation. Its predictions serve as our primary source of semantic observations.
@@
To optimize the semantic 3DGS, we employ a combination of Hellinger distance and cosine similarity losses. Predictions from OneFormer serve as the pseudo-ground truth $\mathcal{P}_{\text{GT}}$, while the rendered semantic outputs are treated as predictions $\mathcal{P}_{\text{pred}}$. To filter out uncertain supervision, we apply an entropy-based mask $M_{H}=\mathbb{I}(H(\mathbf{p})<\tau)$, where $\mathbb{I}(\cdot)$ is the indicator function and $\tau$ is the entropy of a uniform distribution over $k$ categories, i.e. $\tau=\log(k)$. The entropy at each pixel is computed as:
@@
Following SplaTAM [51], our Gaussian Mapping backbone optimizes the map using a subset of keyframes instead of all input frames. Every fifth frame is considered a keyframe candidate, and the map is updated using local keyframes with the highest 3D overlap, computed by backprojecting depth maps and evaluating visibility within keyframe frustums. This provides efficient multiview supervision but may overfit occluded regions, reducing opacity for valid Gaussians behind surfaces.

### v2

@@ 1 Introduction
Mobile robots are expected to play a significant role in human-centered environments, such as warehouses, factories, hospitals, and homes, as well as in dangerous settings, such as mines and nuclear facilities. Rich and accurate geometric and semantic representations are necessary in these scenarios so that robots can understand, interpret, and interact meaningfully with their surroundings. For instance, in automated warehouses, robots are required to recognize various items and place them in the correct sorting zones accordingly. Scene understanding is enabled by a semantic map that is linked to the geometric map [raychaudhuri2025semantic], which represents the spatial layout of an environment, and enriches it with high-level information such as object categories, surface labels, and functional affordances [ruiz2017building, mascaro2022volumetric, mccormac2018fusion, li2025hier]. Such maps are critical for a range of tasks including navigation, inspection, object manipulation, human-robot interaction, and long-term autonomy.
@@
Despite substantial advances in semantic mapping, most current approaches are unable to determine the most informative path for the robot. Instead, they passively rely on externally determined trajectories or predefined exploration strategies [rosinol2020kimera, li2024sgs, ji2024neds, yang2025opengs], leading to incomplete or suboptimal scene understanding. In this paper, we present an approach of active semantic mapping that seeks to close the loop between perception and action. This allows agents to plan their next moves and observations to improve the quality, completeness, and efficiency of the semantic map. Our approach, named ActiveSGM (Active Semantic Gaussian Mapping), is the first active semantic mapping system based on radiance fields, enabling rapid exploration, efficient environment understanding, and high-fidelity real-time rendering, ultimately leading to more intelligent and efficient robotic behaviors. ActiveSGM aims to infer the semantic labels of all visible surfaces, without favoring any particular label.
@@
To select the most informative views for the robot, we seek to quantify both geometric and semantic uncertainty. At the geometric level, uncertainty is typically measured by the expected error in the estimated 3D coordinates [cadena2016past, lluvia2021active, placed2023survey]. At the semantic level, uncertainty estimation primarily captures ambiguity among semantic classes. Recent surveys on semantic uncertainty quantification [matez2024voxeland, wilson2024convbki, wilson2024modeling] found that it is inherently dependent upon the choice of semantic representation.
@@
The choice of semantic representation plays a critical role in semantic mapping systems, which commonly adopt two forms: probability distributions or embeddings. For distribution-based representations, existing methods (e.g., [ronneberger2015u, jain2023oneformer, hu2017maskrnn]) employ either hard or soft assignment strategies. Hard assignments, such as one-hot encoding, strictly assign a single label to each pixel. In contrast, soft assignments allocate a complete categorical probability distribution to each 3D primitive, naturally capturing uncertainty but also incurring higher memory requirements as the number of categories grows. Alternatively, embedding-based representations can also be viewed as a form of soft assignment. Methods like  [yang2025opengs, qin2024langsplat] utilize features, such as those from DINO [caron2021dino] or CLIP [cherti2023reproducible], to encode semantic embeddings. However, these embeddings are high-dimensional, posing challenges for storage and real-time rendering in large scenes. Consequently, some approaches compress the features into lower-dimensional spaces, such as the three-dimensional RGB color space. The dimension of the embedding feature space directly determines the effectiveness of class discrimination. High-dimensional embeddings, like those from DINO or CLIP, provide a sufficiently expressive feature space to effectively distinguish categories. However, as embeddings become compressed, for instance into RGB space, color blending during multi-view reconstruction inevitably occurs, producing blended colors that may correspond to unrelated categories instead of the original ones.
@@
To summarize, we propose the first dense active semantic mapping system built upon a 3D Gaussian Splatting (3DGS) backbone, which integrates semantic-aware mapping and planning for active reconstruction. This enables the robot to construct a more accurate geometric map and a richer semantic map with fewer observations. Our method addresses several key challenges:
@@ 2 Related Work
In this section, we review prior work, starting from dense SLAM, active mapping, semantic mapping, and concluding with active semantic mapping. We focus on methods utilizing either Neural Radiance Fields (NeRF) [mildenhall2020nerf, tewari2022advances, xie2022neural] or Gaussian Splatting (GS) [kerbl20233dgs, fei2024survey, zhu2024_3d] as the representation.
@@
Autonomous robotics relies on foundational capabilities such as localization, mapping, planning, and motion control [siegwart2011introduction]. The need to realize these capabilities has spurred advancements in various areas, including visual odometry [scaramuzza2011visual, zhan2020visual], structure-from-motion (SfM) [schonberger2016structure], and Simultaneous Localization and Mapping (SLAM) [thrun2002probabilistic, durrant2006simultaneous, davison2007monoslam, cadena2016past, campos2021orb]. For surveys of the impact of radiance fields in SLAM and robotics in general, we refer readers to  [bao2025_3dgs_survey, tosi2024nerfs, wang2024nerfrobotics, zhu2024_3d]. Progress in radiance fields has given rise to a multitude of dense SLAM methods, that estimate depth for almost every pixel of the input images, using NeRF (or other implicit representations, such as TSDF) [chung2023orbeez, deng2024plgslam, johari2023eslam, li2023end, stier2023finerecon, sucar2021imap, wang2023coslam, zhang2023go-slam, zhu2022nice, zhu2023nicer] or GS [deng2024compact, huang2024photo, keetha2024splatam, matsuki2024gaussian, yan2024gs-slam, yugay2023gaussian] to represent scene geometry and appearance. We use SplaTAM [keetha2024splatam] as the SLAM backbone of our algorithm.
@@
The goal of SLAM is to estimate the camera/vehicle trajectory from sensor data. Active mapping, or exploration, is a related problem in the domain of active perception [aloimonos1988active, bajcsy1988active], where the goal is guiding the sensor to acquire images beneficial to a downstream task. The most common objectives are to reduce uncertainty, equivalently to increase information gain, [stachniss2009robotic] or to detect and visit frontiers [yamauchi1997frontier]. Prior approaches such as [yamauchi1997frontier, georgakis2022uncertainty, siming2024active] adopt a long-term planning paradigm, in which the information gain is estimated along multiple candidate trajectories before execution. Subsequent studies [feng2024naruto, ramakrishnan2020occupancy, chen2025activegamer], on the other hand, reformulate the problem within a next-best-view framework, where the exploration process is guided by sequential decisions that progressively construct the overall trajectory. Early work demonstrated the effectiveness of active mapping [bourgault2002information, davison2002simultaneous, feder1999adaptive, stachniss2004exploration, thrun1991active], while overviews of the state of the art can be found in [cadena2016past, lluvia2021active, placed2023survey].
@@
Recently, NeRF-based approaches have been applied to path planning [adamkiewicz2022vision] and next-best-view selection [ran2023neurar, lee2022uncertainty, pan2022activenerf], though they are often limited by their high computational cost [kuang2024active]. To overcome these limitations, hybrid models such as ActiveRMAP [zhan2022activermap] integrate implicit and explicit representations. NARUTO [feng2024naruto] introduces an active neural mapping system with 6DoF movement in unrestricted spaces, while Kuang et al. [kuang2024active] integrate Voronoi planning to scale exploration to larger environments. 3DGS offers a faster alternative, making real-time mapping and exploration more feasible. Recent works like ActiveSplat [li2024activesplat] utilize a hybrid map with topological abstractions for efficient planning, ActiveGS [jin2025activegs] also uses a hybrid map and associates a confidence with each Gaussian to guide exploration, and AG-SLAM [jiang2024agslam] incorporates 3DGS with Fisher Information to balance exploration and localization in complex environments. ActiveGAMER [chen2025activegamer] introduces a rendering-based information gain criterion that selects the next-best view for enhancing geometric and photometric reconstruction accuracy in complex environments. RT-GuIDE [tao2024rt] uses a simple uncertainty measure to achieve real-time planning and exploration on a robot. Recently, NextBestPath [li2025nextbestpath] considers longer horizons than just the single next view. Like all the methods in this paragraph, however, it does not consider semantics.
@@
The goal of semantic mapping is to infer scene descriptions that go beyond geometry [raychaudhuri2025semantic]. In general, methods in this category endow their 3D representation with semantic labels, which are inferred via semantic segmentation of the input RGB or RGB-D images. Early work includes approaches such SemanticFusion [mccormac2017semanticfusion], Fusion++ [mccormac2018fusion], PanopticFusion [narita2019panopticfusion] and Kimera [rosinol2020kimera], which have adopted different representations exploring tradeoffs between precision and efficiency. Radiance field-based methods are surveyed by Nguyen et al. [nguyen2024semantically]. Among them, GSNeRF [chou2024gsnerf] introduces the Semantic Geo-Reasoning and Depth-Guided Visual modules to train a NeRF that encodes semantics along with appearance. Wilson et al. [wilson2024modeling] use the variance of the semantic representation at each Gaussian as a proxy for semantic uncertainty. HUGS [zhou2024hugs] jointly optimizes geometry, appearance, semantics, and motion using a combination of static and dynamic 3D Gaussians. Logits for all classes are stored with the Gaussians, but the number of classes is small. All of these approaches operate on all frames in batch mode, however.
@@
Gaussian splats are well suited for semantic mapping because they can encode additional attributes and are amenable to continual learning, unlike NeRF [cai2023clnerf]. All methods below operate on RGB-D video inputs. We point out the important representation choices made by their authors. SGS-SLAM [li2024sgs] augments a GS-based SLAM system with additional test-time supervision via 2D semantic maps. The authors argue that any off-the-self semantic segmentation algorithm can be integrated in SGS-SLAM and use ground truth labels to supervise the splats for simplicity. High-dimensional semantic labels are converted to "semantic colors" to save space. NIDS-SLAM [haghighi2023nids] uses a 2D transformer [cheng2022masked] to estimate keyframe semantics, also converting the semantic labels into "semantic colors". NEDS-SLAM [ji2024neds] reduces the memory footprint of the high-dimensional semantic features obtained by DINO [caron2021dino] to three values per splat via a lightweight encoder. OpenGS-SLAM [yang2025opengs] infers consistent labels via the consensus of 2D foundational models across multiple views. It can handle an open vocabulary, but stores only one label per splat.
To overcome the limited dimensionality of colormaps, researchers have endowed the splats with embeddings of the high-dimensional vectors of logits. SNI-SLAM [zhu2024sni] model the correlations among appearance, geometry and semantic features through a cross-attention mechanism and use feature planes [johari2023eslam] to save memory. DNS-SLAM [li2024dns] relies on a multi-resolution hash-based feature grid. Optimization is performed in latent space, while ground truth 2D semantic maps are used as inputs. SemGauss-SLAM [zhu2024semgauss] augments the splats with a 16-channel semantic embedding and presents semantic-informed bundle adjustment. The paper includes results using ground truth 2D labels for supervision, as well as labels inferred by a classifier operating on DINOv2 [oquab2023dinov2] features. Hier-SLAM [li2025hier] addresses the increased storage requirements via a hierarchical tree representation, generated by a large language model. It can handle over 500 semantic classes, but it is also provided the ground truth semantic maps of the images during optimization.
@@
The above semantic mapping and SLAM approaches are "passive" in the sense that the camera is not actively controlled but follows a predetermined trajectory. In contrast, methods such as [siming2024active, asgharivaskasi2023semantic] actively plan trajectories that maximize the mutual info rmation between past and future semantic observations. Other relevant works have focused on object search using semantic contextual priors—e.g., leveraging knowledge that cups are typically found in kitchens—without explicitly predicting semantic labels for every point in the map [gervet2023navigating, georgakis2022learning, ramakrishnan2022poni]. Among these approaches, more relevant to ours is the work of Zhang et al. [zhang2024active] that relies on semantic mutual information and properties of the SLAM pose graph for metric-semantic active mapping. An octree is used to maintain the map, but the current implementation is limited to 2D motion and 8 classes, while ground truth labels are used as semantic observations. Marza et al. [marza2024autonerf] added a semantic head to Nerfacto [tancik2023nerfstudio] and used it for active mapping of appearance, geometry and semantics. They compared using ground truth semantic labels and Mask-R-CNN [he2017maskrcnn] to detect 15 object categories, and observed large differences in the metrics. Exploration policies are trained using reinforcement learning and consider the 15 object categories. Unlike our approach, the trajectory is restricted to the ground plane. It is not clear how this approach would have to be modified if all semantic classes in the scene would have to be considered.
@@ 3 Method
Figure 1: Overview of the ActiveSGM System. Our framework integrates observation, mapping, and planning into a unified active semantic mapping system. At each time step, posed RGB-D frames along with semantic predictions from OneFormer [jain2023oneformer] are stored in a keyframe database. Selected frames are used to update a Semantic Gaussian Map that encodes geometric, photometric, and semantic properties and is optimized through differentiable rendering. An occupancy-based Exploration Map is updated using the current view and used to sample candidate viewpoints in free space. Next-best views are selected by jointly evaluating geometric and semantic exploration criteria (E.C.), and a path planner navigates toward the selected pose. This closed-loop system enables efficient, semantics-aware reconstruction and exploration in complex 3D environments.
@@ 3.1 Semantic Gaussian Mapping
Gaussian Mapping leverages 3DGS to represent scenes as collections of 3D Gaussians, encoding both appearance and geometry for real-time rendering of high-fidelity color and depth images. Building upon the work of Kerbl et al. [kerbl20233dgs], we adopt the streamlined approach proposed in SplaTAM [keetha2024splatam]. This method employs isotropic Gaussians with view-independent color, optimizing parameters such as color (c), center position ($\boldsymbol{\mu}$), radius ($r$), and opacity ($o$). A notable advantage of 3DGS is its capability for real-time rendering, enabling the synthesis of high-fidelity color and depth images from arbitrary camera poses. This is achieved by transforming 3D Gaussians into camera space, sorting them front-to-back, projecting them onto the 2D image plane, and employing alpha-blending for compositing. The color, depth, and silhouette at pixel $\mathbf{p}$ are rendered from the Gaussian map, where the silhouette indicates whether $\mathbf{p}$ receives a significant projection from any Gaussian. The general rendering process is formulated as
@@
We incorporate semantics into the 3DGS map by using OneFormer [jain2023oneformer], a state-of-the-art model for unified segmentation, to perform semantic segmentation. Its predictions serve as our primary source of semantic observations.
@@
To optimize the semantic 3DGS, we employ a combination of the Hellinger distance and cosine similarity losses. Predictions from OneFormer serve as the pseudo-ground truth $\mathcal{P}_{\text{GT}}$, while the rendered semantic outputs are treated as predictions $\mathcal{P}_{\text{pred}}$. To filter out uncertain supervision, we apply an entropy-based mask $M_{H}=\mathbb{I}(H(\mathbf{p})<\tau)$ on $\mathcal{P}_{\text{GT}}$, where $\mathbb{I}(\cdot)$ is the indicator function and $\tau$ is the entropy of a uniform distribution over $k$ categories, i.e. $\tau=\log(k)$. The entropy at each pixel is computed as:
@@
Following SplaTAM [keetha2024splatam], our Gaussian Mapping backbone optimizes the map using a subset of keyframes instead of all input frames. Every fifth frame is considered a keyframe candidate, and the map is updated using local keyframes with the highest 3D overlap, computed by backprojecting depth maps and evaluating visibility within keyframe frustums. This provides efficient multiview supervision but may overfit occluded regions, reducing opacity for valid Gaussians behind surfaces.